# Modelling reliability of reversible circuits with 2D second-order cellular automata

### v2

@@ Abstract
journal: arXiv{graphicalabstract}
@@ 1 Introduction
Reversible computations have quite a long history Ben73 and nowadays they deserve increasing attention due to different applications from the reduction of power consumption to the theory of quantum computations KM17 ; COST20 ; RC22 ; RC23 . The analysis of uncertainties for the testing of usual logic circuits is well developed area with variety of different methods Circ13 , but some subtleties can appear for reversible circuits. The problems of faults and the reliability of the reversible circuits were also covered quite widely PHM03 ; HPB04 ; RTest20 .
From the one side the testing of reversible circuits may be even simpler for some particular faults PHM03 , but some difficulties may appear in the more general cases HPB04 . There are also some problems due to specific properties of reversible circuits. From the one hand, the reversibility produces an advantage due to negligible power consumption, from the other one, the undesirable perturbations may spread without dissipation and produce the problems with too high error sensitivity.
More detailed analysis of the particular structures used for construction of the reversible logical circuits can be useful for the sensitivity analysis and consideration of some other problems. A model of so-called quantum-dot cellular automata can be used as some example RTest20 ; QDCA07 . The presented paper is devoted to the more abstract models based on the reversible cellular automata (RCA).
The consideration of the uncertainty propagation can be described with irreversible probabilistic cellular automata UQCA , however an application of similar ideas to the reversible case produces some problems if to try add the noise directly into update rule of the cells, because such a system could not perform the reliable computations RCAnoise .
The issue with noisy local rules of the probabilistic RCA justifies sensitivity analysis of usual (deterministic) RCA evolution with respect to different perturbations of the initial configurations. The family of second-order RCA with four states considered in this work uses the standard method of construction from the cellular automata with two states M84 ; V84 ; TM90 .
The presented work is in some agreement with an earlier suggestion from Ref. TM90 ‘[…] the most practical approach is to start with a local map that directly supports logic gates and wires, and then build the appropriate logic circuits out of these primitives’. Examples of such approach often inspired by so-called billiard ball model (BBM) KM17 ; FT82 ; JK18 . A computationally universal second-order RCA derived from cellular automaton with three states was discussed already in Ref. M84 .
Another approach to the computational universality in RCA uses 1D design based on universal Turing machine and it is discussed together with 2D logic circuits based on so-called partitioned CA (PCA) in Ref. KM17 ; JK18 . PCA is usual choice for modelling of reversible circuits due to natural relation with BBM.
@@
Models without conservation of cell number may be necessary in such a case. In irreversible CA undesired fast growth of number of cells could be corrected by modification of local rules to cancel some configurations. For example, diverse evolution Game of Life CA WW ; Gard and possibility of universal computations with such a model supported by rules with good balance between appearance and cancellation of different configurations.
However, there is specific problem with RCA, because reversible evolution make impossible extinction of any configuration and without conservation of cell number (or some other non-negative quantity) undesired growth and chaotic behaviour of initially regular configuration is rather common result of RCA evolution in unbounded domain.
Few reversible second-order RCA discovered by author some time ago and discussed in this work provide interesting combination of stable and expansive behaviour. From the one hand such RCA can be used for creation of reversible universal gates, from the other one even expansive behaviour displays some regularities and fractal patterns. Investigation sensitivity and reliability of reversible circuits with such models can be quite instructive and relevant to analogue processes in more realistic systems.
Let us briefly outline contents of this paper. After recollection of basic properties and possible classification of CA in Sec. 2.1 and Sec. 2.2 few original second-order RCA for circuits construction are introduced in Sec. 2.3 and examples of logic gates for universal computations are described in Sec. 2.4. Different kinds of errors from the defects and the improper signal interactions are discussed in Sec. 3 recollecting some quantities such as Lyapunov exponents and formal dimensions those could be useful for analysis of the damage spread together with the numerical experiments. In conclusion, Sec. 4 summarizes this work.
@@ 2.1 Basic properties and classification of CA
Let us recollect some properties of cellular automata (CA) on the rectangular grids with dimension $d$ often denoted as $\mathbb{Z}^{d}$ WolCA ; JK05 . Two-dimensional CA are quite natural for modelling of the logic circuits. The state of any cell on the grid $\mathbb{Z}^{d}$ belongs to the finite-dimensional set $S$ with $k$ elements and for simplicity may be described by a number from interval $10,\ldots,k-$ 1. State $k=0$ could be considered as an empty cell and there are $k-1$ kinds of occupied cells. In the simplest case with two states the cell is either occupied or not.
Evolution of such a system is described by synchronous update at discrete time steps. The local update rule for each cell $c_{\mathbf{z}}$, $\mathbf{z}\in\mathbb{Z}^{d}$ is a function depending only on few cells in the certain neighborhood described by $n$ vectors of relative displacements $\mathcal{N}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{n})$
@@
The reversible cellular automata (RCA) are appropriate for the purposes of presented work, but there are no effective algorithms for inverting given local rule for $d>1$ JK18 . However, some methods allow to construct new local rules with known simple inverse from the very beginning.
One of these methods uses so-called second-order CA M84 ; V84 ; TM90 ; JK18 ; Ila recollected below. The second-order CA uses values from two previous steps for calculating of a state for a next one
@@
Let us now consider arbitrary CA with local map $f$ Eq. (1). It is possible to define new second-order CA
@@
where $\mathbin{\ominus_{k}}$ is subtraction modulo $k$ and such CA is reversible M84 ; V84 ; TM90 ; Ila with the inverse map
@@
For an initial CA with two states $k=2$ the operation $\mathbin{\ominus_{2}}$ coincides with XOR (eXclusive OR often denoted as $\oplus$) producing more familiar examples JK18 . The second-order RCA can be considered as first-order RCA with state spaces $S\times S$ and local rule TM90 ; JK18
@@ 2.2 Classification of CA and RCA
Due to Eq. (5) there is one to one correspondence between the second-order RCA with four states and two-state CA. Thus, classification of 2D CA with two states can be also used for such RCA. A natural choice is the isotropic local rules also known as completely symmetric PW85 i.e., invariant with respect to rotations and reflections. However, even with such restriction there are $2^{102}\simeq 5\times 10^{30}$ different CA PW85 .
There are more narrow classes of CA, for example totalistic with value $f$ in Eq. (1) is depending only on sum of occupied cells in neighborhood and outer totalistic with taking into account the sum and the value of cell itself PW85 . The last example is also known as semi-totalistic or Life-like CA (GoL10, , Ch. 6) due to famous Conway’s Game of Life CA WW ; Gard and there are $2^{18}\simeq 3\times 10^{5}$ such rules PW85 .
For the certainty the term outer totalistic is saved further for particular case of outer totalistic inner-independent CA (GoL10, , Ch. 13) with local rule taking into account only ‘outer neighborhood’ without cell itself.
Between isotropic and totalistic classes there is an intermediate one with separate counts for two kinds of neighboring cells: the four closest neighbors have a common edge with central cell and another four cells have only common corner. Such a class is also known as corner-edge totalistic (CET) or quarter-totalistic CA rules term and it is convenient again to consider inner-independent case without dependence on central cell denoted further as corner-edge outer totalistic (CEOT) CA.
@@
The similar method can be used to calculate number of possible local rules in other classes of CA with two states discussed above and corresponding second-order RCA. The considered classes also can be convenient due to relations with invariant quantities in second-order RCA defined by count of cell values V84 ; Ila ; YP84 ; GV86 ; ST12 . However, there are few alternative methods of classifications class and these topics should be discussed elsewhere.
@@ 2.3 RCA family for circuits construction
Let us consider subfamily of CEOT CA with local update rules producing occupied cell only if all cells with common corner are empty, i.e., only pairs with $n_{c}=0$ may appear in the sequence for $f(p)=1$ discussed above. Selected CA are between simplest examples of CA from considered families (CET, CEOT) and were found by not very extensive search for CA with universal gates. Anyway, analogue CA were not found in similar way between totalistic rules and so CEOT was used instead. The simple example of such CA denoted here as $C_{0}E_{1}$ includes single pair with $n_{c}=0$, $n_{e}=1$ and second-order RCA derived from such a rule using Eq. (5) is denoted as $C_{0}E_{1}^{\,I\!I}$. Few different extensions of such RCA defined further could be denoted as $C_{0}E_{1\dots}^{\,I\!I}$.
Condition about empty cells with common corner is quite restrictive and many configurations of initial irreversible CA disappear after single step, but in Eq. (5) for second-order automata $f(p)=0$ is applied as XOR operation to the previous state. Configurations corresponding to those that would disappear in the original CA with two states become stable in the second-order RCA if all cells have the state $s=3$ with two units in binary notation or blink between two values for $s=1,2$.
@@ 2.4 Constructions of reversible gates
An analogy of BBM computations FT82 can be developed with shortest three-cell signals (‘ants’). The simplest encoding is supposed in such a case with the unit and zero are represented by presence or absence of an elementary signal respectively. It has some disadvantages, but demonstrates many basic principles of circuits design and more difficult encodings should be discussed elsewhere.
@@
The important element for BBM is the reversible Fredkin gate FT82 . Such a gate (with three inputs and three outputs) performs conditional exchange between two data signals depending on value of a control one. The action of Fredkin gate can be formally expressed as
@@
Here two data signals are exchanged if control signal presents. Such convention formally differs from initially suggested in Ref. FT82 , but it can be simply altered by exchange of data outputs.
@@
The Fredkin gates together with delay elements, swaps and ‘constants’ (i.e. some amounts of auxiliary signals) allow to construct any logic circuits FT82 . Necessity of some constants for used encoding is clear even from construction of NOT gate that can be expressed with Fredkin gate Eq. (6) as
@@
Similarly with BBM construction of the Fredkin gate can be expressed with simpler switch gates FT82 for conditional routing of one data signal by a control one. The configuration of the switch gate for discussed RCA is shown on the left side of Figure 2. The scheme of the switch gate shown on the right side of the figure has been modified compared to that used in FT82 to more accurately display the directions of outgoing signals.
@@
The scheme of Fredkin gate composed from four such switch gates initially proposed by R. Feynman and A. Ressler may be found in Ref. FT82 . Possible realization for considered RCA is shown on Figure 3. The simplified circuit diagram without delay elements shown at the bottom of Figure 3 uses two switch gates together with two inverted switch gates. All switch gates on Figure 3 are flipped vertically compared with Figure 2 and surrounded by dashed rectangles on the diagram.
@@
where $\oplus$ is addition modulo 2 (XOR). Such gate can be represented as
@@
Possible design is shown on Figure 5. The simplified circuit diagram without delay elements shown at the bottom of Figure 5 uses a gate depicted on Figure 4 (surrounded by a dashed rectangle) together with inversion of such gate and exchanges paths of two signals between them.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ was briefly mentioned first in Ref. AV14 , $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ were introduced and implemented in software developed by author with different names around the same time CART together with yet another isotropic RCA.
@@
It was already mentioned, that the second-order RCA can be considered as usual (first-order) RCA with four states. Using conversion into appropriate format for local rules with four states some numerical experiments discussed here was also performed on general software for CA simulation developed by A. Trevorrow, T. Rokicki et al GoL .
@@ 3.1 Damage propagation from single defect
Quantification of damage propagation in CA can be useful for reliability analysis of logic circuits based on such a model. Informal idea of Lyapunov exponents sometimes used as a measure of such propagation WolCA and few approaches to rigorous definition are developed BBR92 ; MS92 ; Ti00 ; BB10 ; BG18 ; Ko21 despite of certain difficulties and ambiguities related with application of such quantity to discrete systems. An intensity of defect accumulation and speed of the damage spread could be mentioned as two alternative views on the Lyapunov exponent BG18 .
@@
Such a problem could be addressed by consideration of separate replica of a system for each defect for any time step BBR92 with formal use of so-called Boolean derivatives Vi90 , but such calculation without proper accounting of defects cancellation may produce overestimation of possible effects.
@@
where $\oplus$ is addition modulo 2. According to enumeration in Ref. W93 the CA corresponds to rule 90 and provides important example of additive CA AlgCA . Statistical properties of second-order RCA produced by such a rule denoted as 90R was discussed in Ref. Ta87 .
For initial (irreversible) rule 90 CA calculation with Boolean derivatives produces value $\lambda=\ln 2$ for Lyapunov exponent BBR92 . For configuration with single occupied cell the total number of non-empty cells after $n$ steps is described by sequence A001316 in OEIS OEIS and can be expressed as $2^{wt({n})}$, where $wt({n})$ is number of units in binary expansion of $n$ (Hamming weight).
Despite of exponential notation the value $2^{wt({n})}$ may not exceed $n+1$, but for any even $n$ the value doubles on next step $n+1$ resembling dynamics that could be expected for process with $\lambda=\ln 2$. Damage front for such CA moves with unit speed in both directions and could provide only minimal information for comparison with other CA BBR92 .
@@
This CA is notable as a simple example with self-replication of any pattern discovered by E. Fredkin about 1960 and recollected by M. Gardner (Gard, , Ch. 21). For a single occupied cell after $n$ steps there are $4^{wt({n})}$ non-empty cells, see sequence A102376 in OEIS. For particular values of $t$ the configurations correspond to the square grids with different number of points and intervals between them. Due to Eq. (10) the CA is linear with respect to addition modulo 2 and because any pattern can be formally represented as a sum of few configurations with single occupied cell, at appropriate time steps the pattern is replicated over grids with sufficiently large intervals.
Using earlier conventions for CEOT rules such CA could be formally denoted as $C_{0-4}E_{1,3}$, but the notation $E_{1,3}$ is more natural for this CA with von Neumann neighborhood, because the local rule Eq. (10) does not depend on number of cells with common corner. The CA should be distinguished from similar replicator with Moore neighborhood SlCA , see sequence A160239 in OEIS.
Second-order 2D RCA from $E_{1,3}$ can be denoted as $E_{1,3}^{I\!I}$. Number of cells with different states after $n$ steps for configuration with a single cell with state 1 (red) is described by sequences A244642 and A244643 in OEIS. The RCA $E_{1,3}^{I\!I}$ is not replicator, but it may produce interesting fractal-like patterns. For configuration with single red cell evolution is also coincides with considered earlier RCA $C_{0}E_{1}^{\,I\!I}$ AV14 and for $C_{0}E_{1,3}^{\,I\!I}$ such coincidence also can be shown in a similar way due to some similarity of local rules with $C_{0}E_{1}^{\,I\!I}$ and $E_{1,3}^{I\!I}$, see Figure 6.
@@
The numbers $N_{1}$ and $N_{2}$ of cells with states 1 and 2 after $n$ steps for configuration with single red cell satisfy recursive formulas AV14
@@
The first part of each equation would correspond to asymptotically quadratic growth $O(n^{2})$, but the second part to only linear one $O(n)$. Such a variance in formal estimation of dimensions seems to be in some agreement with the shape of configurations Figure 6 resembling some geometrical sequences used for construction of fractals Frac .
Such behaviour justifies consideration of models with polynomial growth together with an analogy of formal or fractal dimension Frac already investigated earlier in context of CA (Ila, , Ch. 4) and denoted further as ${\mathfrak{D}}$ that is not vanishing unlike $\lambda$ Eq. (8)
@@
The subtleties of transition from definition of fractal dimension relevant to continuous case to such discrete systems as CA can be found in Ref (Ila, , Ch. 4), but here formal dimension ${\mathfrak{D}}$ is treated rather as convenient quantity defined by Eq. (12).

### v3

@@ Abstract
journal: Physica D (Nonlinear Phenomena)
{graphicalabstract}[Uncaptioned image]
@@ 1 Introduction
The reversible computations have quite a long history [1] and nowadays they deserve an increasing attention due to different applications from the reduction of power consumption to the theory of quantum computations [2, 3, 4, 5]. The analysis of uncertainties for the testing of usual logic circuits is well developed area with variety of different methods [6], but some subtleties can appear for the reversible circuits. The problems of faults and the reliability of the reversible circuits were also covered quite widely [7, 8, 9].
Sometimes the testing of reversible circuits may be even simpler for some particular faults [7], but some difficulties may appear in the more general cases [8]. There are also some problems due to specific properties of the reversible circuits. From the one hand, the reversibility produces an advantage due to negligible power consumption, from the other one, the undesirable perturbations may spread without dissipation and produce the problems with too high error sensitivity.
More detailed analysis of the particular structures used for construction of the reversible logic circuits can be useful for the sensitivity analysis and consideration of some other problems. A model of so-called quantum-dot cellular automata can be used as some example [9, 10]. The presented paper is devoted to the more abstract models based on the reversible cellular automata (RCA).
The consideration of the uncertainty propagation can be described with irreversible probabilistic cellular automata [11], however an application of similar ideas to the reversible case produces some problems if to try add the noise directly into update rule of the cells, because such a system could not perform the reliable computations [12].
The issue with noisy local rules of the probabilistic RCA justifies sensitivity analysis of usual (deterministic) RCA evolution with respect to different perturbations of the initial configurations. The family of second-order RCA with four states considered in this work uses the standard method of construction from the cellular automata with two states [13, 14, 15].
The presented work is in some agreement with an earlier suggestion from Ref. [15] ‘[…] the most practical approach is to start with a local map that directly supports logic gates and wires, and then build the appropriate logic circuits out of these primitives’. Examples of such approach often inspired by so-called billiard ball model (BBM) [2, 16, 17]. A computationally universal second-order RCA derived from cellular automaton with three states was discussed already in Ref. [13].
Another approach to the computational universality in RCA uses 1D design based on universal Turing machine and it is discussed together with 2D logic circuits based on so-called partitioned CA (PCA) in Ref. [2, 17]. PCA is usual choice for modelling of reversible circuits due to natural relation with BBM.
@@
Models without conservation of cell number may be necessary in such a case. In irreversible CA undesirable rapid growth of number of cells could be corrected by modification of local rules to cancel some configurations. For example, diverse evolution Game of Life CA [20, 21] and possibility of universal computations with such a model is supported by rules with good balance between appearance and cancellation of different configurations.
However, there is specific problem with RCA, because reversible evolution makes impossible extinction of any configuration and without conservation of cell number (or some other non-negative quantity) undesirable growth and chaotic behaviour of initially regular configuration is rather common result of RCA evolution in unbounded domains.
A few reversible second-order RCAs discovered by author some time ago and discussed in this work provide interesting combination of stable and expansive behaviour. From the one hand such RCA can be used for creation of reversible universal gates, from the other one even expansive behaviour displays some regularities and fractal patterns. Investigation sensitivity and reliability of reversible circuits with such models can be quite instructive and relevant to analogue processes in more realistic systems.
Let us briefly outline contents of this paper. After recollection of basic properties and possible classification of CA in Sec. 2.1 and Sec. 2.2 a few original second-order RCAs for circuits construction are introduced in Sec. 2.3 and examples of logic gates for universal computations are described in Sec. 2.4. Different kinds of errors from the defects and the improper signals interactions are discussed in Sec. 3 recollecting some quantities such as Lyapunov exponents and formal dimensions those could be useful for analysis of the damage spread together with the numerical experiments. In conclusion, Sec. 4 summarizes this work.
@@ 2.1 Basic properties and classification of CA
Let us recollect some properties of cellular automata (CA) on the rectangular grids with dimension $d$ often denoted as $\mathbb{Z}^{d}$ [22, 23]. Two-dimensional CA are quite natural for modelling of the logic circuits. The state of any cell on the grid $\mathbb{Z}^{d}$ belongs to the finite-dimensional set $S$ with $k$ elements and for simplicity may be described by a number from interval $10,\ldots,k-$ 1. State $k=0$ could be considered as an empty cell and there are $k-1$ kinds of occupied cells. In the simplest case with two states the cell is either occupied or not.
Evolution of such a system is described by synchronous update at discrete time steps. The local update rule for each cell $c_{\mathbf{z}}$, $\mathbf{z}\in\mathbb{Z}^{d}$ is a function dependent on only a few cells in the certain neighborhood described by $n$ vectors of relative displacements $\mathcal{N}=(\mathbf{v}_{1},\ldots,\mathbf{v}_{n})$
@@
The reversible cellular automata (RCA) are appropriate for the purposes of presented work, but there are no effective algorithms for inverting given local rule for $d>1$ [17]. However, some methods allow to construct new local rules with known simple inverse from the very beginning.
One of these methods uses so-called second-order CA [13, 14, 15, 17, 18] recollected below. The second-order CA uses values from two previous steps for calculating of a state for a next one
@@
Let us now consider arbitrary CA with a local update rule $f$ Eq. (1). It is possible to define new second-order CA
@@
where $\mathbin{\ominus_{k}}$ is subtraction modulo $k$ and such CA is reversible [13, 14, 15, 18] with the inverse rule
@@
For an initial CA with two states $k=2$ the operation $\mathbin{\ominus_{2}}$ coincides with XOR (eXclusive OR often denoted as $\oplus$) producing more familiar examples [17]. The second-order RCA can be considered as first-order RCA with state spaces $S\times S$ and local rule [15, 17]
@@ 2.2 Classification of CA and RCA
Due to Eq. (5) there is unique relation between the second-order RCA with four states and two-state CA. Thus, classification of 2D CA with two states can be also used for such RCA. A natural choice is the isotropic local rules also known as completely symmetric [24], i.e., invariant with respect to rotations and reflections. However, even with such restriction there are $2^{102}\simeq 5\times 10^{30}$ different CA [24].
There are more narrow classes of CA, for example totalistic with value $f$ in Eq. (1) is depending only on sum of occupied cells in neighborhood and outer totalistic with taking into account the sum and the cell itself [24]. The last example is also known as semi-totalistic or Life-like CA [25, Ch. 6] due to famous Conway’s Game of Life CA [20, 21] and there are $2^{18}\simeq 3\times 10^{5}$ such rules [24].
For the certainty the term outer totalistic is saved further for particular case of outer totalistic inner-independent CA [25, Ch. 13] with local rule taking into account only ‘outer neighborhood’ without cell itself.
Between isotropic and totalistic classes there is an intermediate one with separate counts for two kinds of neighboring cells: the four closest neighbors have a common edge with central cell and another four cells have only common corner. Such a class is also known as corner-edge totalistic (CET) or quarter-totalistic CA rules [26] and it is convenient again to consider inner-independent case without dependence on central cell denoted further as corner-edge outer totalistic (CEOT) CA.
@@
The similar method can be used to calculate number of possible local rules in other classes of CA with two states discussed above and corresponding second-order RCA. The considered classes also can be convenient due to relations with invariant quantities in second-order RCA defined by count of cell values [14, 18, 27, 28, 29]. However, there are few alternative methods of classifications [30] and these topics should be discussed elsewhere.
@@ 2.3 RCA family for circuits construction
Let us consider subfamily of CEOT CA with local update rules producing occupied cell only if all cells with common corner are empty, i.e., only pairs with $n_{c}=0$ may appear in the sequence for $f(p)=1$ discussed above. Selected CA are between simplest examples of CA from considered families (CET, CEOT) and were found by not very extensive search for CA with universal gates. Anyway, analogue CA were not found in similar way between totalistic rules and so CEOT was used instead.
The simple example of such CA denoted here as $C_{0}E_{1}$ has local rule with $f(p)=1$ only for single pair $(n_{c}=0,n_{e}=1)$ and second-order RCA derived from such a rule using Eq. (5) is denoted as $C_{0}E_{1}^{\,I\!I}$. Few different extensions $C_{0}E_{1\dots}^{\,I\!I}$ of such RCA defined further include more pairs with $n_{c}=0$, $n_{e}\geq 1$ in the sequence for $f(p)=1$.
Condition about empty cells with common corner is quite restrictive and many configurations of initial irreversible CA disappear after single step, but in Eq. (5) for second-order automata $f(p)=0$ is applied as XOR operation to the previous state. Configurations corresponding to those that would disappear in the original CA with two states become stable in the second-order RCA if all cells have the state $s=3$ (with two units in binary notation) or blink between two values for $s=1,2$.
@@ 2.4 Constructions of reversible gates
An analogy of BBM computations [16] can be developed with shortest three-cell signals (‘ants’). The simplest encoding is supposed in such a case with the unit and zero are represented by presence or absence of an elementary signal respectively. It has some disadvantages, but demonstrates many basic principles of circuits design and more difficult encodings should be discussed elsewhere.
@@
The important element for BBM is the reversible Fredkin gate [16]. Such a gate (with three inputs and three outputs) performs conditional exchange between two data signals depending on value of a control one. The action of Fredkin gate can be formally expressed as
@@
Here two data signals are exchanged if control signal presents. Such convention formally differs from initially suggested in Ref. [16], but it can be simply altered by exchange of data outputs.
@@
The Fredkin gates together with delay elements, swaps and ‘constants’ (i.e. some amounts of auxiliary signals) allow to construct any logic circuits [16]. Necessity of some constants for used encoding is clear even from construction of NOT gate that can be expressed with Fredkin gate Eq. (6) as
@@
Similarly with BBM construction of the Fredkin gate can be expressed with simpler switch gates [16] for conditional routing of one data signal by a control one. The configuration of the switch gate for discussed RCA is shown on the left side of Figure 2. The scheme of the switch gate shown on the right side of the figure has been modified compared to that used in [16] to more accurately display the directions of outgoing signals.
@@
The scheme of Fredkin gate composed from four such switch gates initially proposed by R. Feynman and A. Ressler may be found in Ref. [16]. Possible realization for considered RCA is shown on Figure 3. The simplified circuit diagram without delay elements shown at the bottom of Figure 3 uses two switch gates together with two inverted switch gates. All switch gates on Figure 3 are flipped vertically compared with Figure 2 and surrounded by dashed rectangles on the diagram.
@@
where $\oplus$ is addition modulo 2 (XOR). Such a gate can be represented as
@@
Possible design is shown on Figure 5. The simplified circuit diagram without delay elements shown at the bottom of Figure 5 uses a gate depicted on Figure 4 (surrounded by a dashed rectangle) together with inversion of this gate and exchanges paths of two signals between them.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ was briefly mentioned first in Ref. [31], $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ were introduced and implemented in software developed by author with different names around the same time [32] together with yet another isotropic RCA.
@@
It was already mentioned, that the second-order RCA can be considered as usual (first-order) RCA with four states. Using conversion into appropriate format for local rules with four states some numerical experiments discussed here was also performed on general software for CA simulation developed by A. Trevorrow, T. Rokicki, et al [33].
@@ 3.1 Damage propagation from single defect
Quantification of damage propagation in CA can be useful for reliability analysis of logic circuits based on such a model. Informal idea of Lyapunov exponents sometimes used as a measure of such propagation [22] and few approaches to rigorous definition are developed [34, 35, 36, 37, 38, 39] despite of certain difficulties and ambiguities related with application of such quantity to discrete systems. An intensity of defect accumulation and speed of the damage spread could be mentioned as two alternative views on the Lyapunov exponent [38].
@@
Such a problem could be addressed by consideration of separate replica of a system for each defect for any time step [34] with formal use of so-called Boolean derivatives [40], but such calculation without proper accounting of defects cancellation may produce overestimation of possible effects.
@@
where $\oplus$ is addition modulo 2. According to enumeration in Ref. [41] the CA corresponds to rule 90 and provides important example of additive CA [42]. Statistical properties of second-order RCA produced by such a rule denoted as 90R was discussed in Ref. [43].
For initial (irreversible) rule 90 CA calculation with Boolean derivatives produces value $\lambda=\ln 2$ for Lyapunov exponent [34]. For configuration with single occupied cell the total number of non-empty cells after $n$ steps is described by sequence A001316 in OEIS [44] and can be expressed as $2^{wt({n})}$, where $wt({n})$ is number of units in binary expansion of $n$ (Hamming weight).
Despite of exponential notation the value $2^{wt({n})}$ may not exceed $n+1$, but for any even $n$ the value doubles on next step $n+1$ resembling dynamics that could be expected for process with $\lambda=\ln 2$. Damage front for such CA moves with unit speed in both directions and could provide only minimal information for comparison with other CA [34].
@@
This CA is notable as a simple example with self-replication of any pattern discovered by E. Fredkin about 1960 and recollected by M. Gardner [21, Ch. 21]. For a single occupied cell after $n$ steps there are $4^{wt({n})}$ non-empty cells, see sequence A102376 in OEIS. For particular values of $t$ the configurations correspond to the square grids with different number of points and intervals between them. Due to Eq. (10) the CA is linear with respect to addition modulo 2 and because any pattern can be formally represented as a sum of few configurations with single occupied cell, at appropriate time steps the pattern is replicated over grids with sufficiently large intervals.
Using earlier conventions for CEOT rules such CA could be formally denoted as $C_{0-4}E_{1,3}$, but the notation $E_{1,3}$ is more natural for this CA with von Neumann neighborhood, because the local rule Eq. (10) does not depend on number of cells with common corner. The CA should be distinguished from similar replicator with Moore neighborhood [45], see sequence A160239 in OEIS.
Second-order 2D RCA from $E_{1,3}$ can be denoted as $E_{1,3}^{I\!I}$. Number of cells with different states after $n$ steps for configuration with a single cell with state 1 (red) is described by sequences A244642 and A244643 in OEIS. The RCA $E_{1,3}^{I\!I}$ is not replicator, but it may produce interesting fractal-like patterns. For configuration with single red cell evolution is also coincides with considered earlier RCA $C_{0}E_{1}^{\,I\!I}$ [31] and for $C_{0}E_{1,3}^{\,I\!I}$ such coincidence also can be shown in a similar way due to some similarity of local rules with $C_{0}E_{1}^{\,I\!I}$ and $E_{1,3}^{I\!I}$, see Figure 6.
@@
The numbers $N_{1}$ and $N_{2}$ of cells with states 1 and 2 after $n$ steps for configuration with single red cell satisfy recursive formulas [31]
@@
The first part of each equation would correspond to asymptotically quadratic growth $O(n^{2})$, but the second part to only linear one $O(n)$. Such a variance in formal estimation of dimensions seems to be in some agreement with the shape of configurations Figure 6 resembling some geometrical sequences used for construction of fractals [46].
Such behaviour justifies consideration of models with polynomial growth together with an analogy of formal or fractal dimension [46] already investigated earlier in context of CA [18, Ch. 4] and denoted further as ${\mathfrak{D}}$ that is not vanishing unlike $\lambda$ Eq. (8)
@@
The subtleties of transition from definition of fractal dimension relevant to continuous case to such discrete systems as CA can be found in Ref [18, Ch. 4], but here formal dimension ${\mathfrak{D}}$ is treated rather as convenient quantity defined by Eq. (12).